OpenStudy (anonymous):

if a bolt of fabric measuring 20 yards 8 inches is divided into 26 equal sections, how long is each section ?

OpenStudy (anonymous):

how many inches in a yard plus 8 and divide that by 26. Simple

OpenStudy (anonymous):

you figure it out yet ?

OpenStudy (anonymous):

1 foot 10 inches

OpenStudy (tkhunny):

You can work in any units you like. 20 yds + 8 in = 20 yds + (2/3) ft = 20 + (2/9) yds 20 + (2/9) yds / 26 = (7/9) yds = 2 + (1/3) ft = 2 ft 4 in OR 20 yds + 8 in = 728 in 728 in / 26 = 28 in = 2 ft 4 in